Given a two-tailed test, α = .05, N = 12 and H0: ρ ≠ 0, which of these is the correct rcv? (You should refer to a table of critical values of r for this question.)

A. ± .576

B. -.497

C. .576

D. ± .497

SSX = 172.5, SSY = 80.6, and Σ[(X – MX) (Y – MY)] = 95.33, r = :

A. -.81

B. 0

C. .08

D. .81

In a study of the relationship between the tempo of preferred music and daily mood, r2 = 33.11 percent. Which of these is the most appropriate statement, given these findings?

A. The tempo of preferred music causes a 33 percent change in a person's daily mood

B. A person's daily mood explains 33 percent of the variability in the tempo of preferred music

C. The tempo of preferred music explains 33 percent of the variability in a person's daily mood

D. The tempo of preferred music has a moderate influence on a person's daily mood

Using the results r(128) = .34, p > .05, what is sr?

A. .18

B. .08

C. .09

D. .01
Answer #1

Table of Critical Values: Pearson Correlation

1-tailed 0.05 0.025 0.005 2-tailed Degrees of Freedom = N-2 N 0.1 0.05 0.01 1 0.988 0.997 0.999 2 0.900 0.950 0.990 0.805 0.8

Note for N=12,and two tailed test, the correct rcv at \alpha =0.05 is ​​​​\pm0.576. (LOOK AT N=10,2 TAILED \alpha =0.05)

A. IS CORRECT OPTION.

r=E[lx-Mx) ( Y-MX)] = 95.33 * [(x-M-) (Y-Hy)] = 15:33 =) cou(X,Y) 95.33 12 SSX = 172.5 Х sty=8006 Ssx V(X) ssy July 172.5 --- 8

In a study of the relationship between the tempo of preferred music and daily mood, r2 = 33.11 percent. Which of these is the most appropriate statement, given these findings?

C is correct option.

Because r^2 is the proportion of variability in dependent variable (mood) attributable to independent variable(tempo of music)

Using the results r(128) = .34, p > .05, what is sr?